Since the last report, much visible progress has been made, as the LS1 programme approaches the halfway point. From early October, technical and safety shift-crew have been present around the clock, allowing detectors to stay switched on overnight, ensuring that safety systems are operational and instructions for non-expert shift-crew are clear.

LS1 progress

Throughout the summer, whilst the solenoid vacuum tank and YB0 surfaces were accessible, an extensive installation programme took place to prepare for Tracker colder operation and the PLT installation, in 2014, the Phase 1 Pixel Tracker installation, in 2016–’17, and the HCAL Phase 1 upgrade completion, ending in LS2. This included pipework for N2 or dry air to flush the Tracker bulkhead region, many sensors to monitor temperature and dew point in the Tracker and its service channels, heating wires outside the Tracker cooling bundles, supports for the new vacuum-jacketed, concentric, CO2 Pixel cooling lines, the PLT cooling line now necessary following the decision to change to silicon sensors and additional power lines to the HCAL RBX’s to accommodate the Phase 1 upgrade.

The logistic configuration then changed to allow priority access to the +z endcap disks and the –z barrel wheels (with the +z barrel wheels thus closed over the vac tank). Repair and maintenance thus took place on the DT and RPC chambers in YB–1 and YB–2, and those accessible from the –z face of YB0. This work included a major gas leak repair campaign for the barrel RPCs. In addition the current HPD photomultipliers of HO were exchanged for SiPMs. The –z wheels are now closed over the vac tank and RPC repairs are being completed on YB–2. As DT maintenance on barrel wheels is completed, the sector collector crates of the DTs are being replaced by optical drivers and moved from the balconies in the UXC into the USC. Open access to the –z nose has allowed ME1/1 patch panels to be replaced and HE source calibration to be prepared.

Concurrently, the YE4 (+z) disk has been assembled on YE+3, moved to the +z end headwall and detached onto newly installed mountings on the blockhouse. The new ME4/2 chambers have recently been mounted and cabled on YE+3; their commissioning is expected to finish near the beginning of the CMS Week (at the time of submission of this report). Meanwhile, installation of the new RE4/2 super-modules, covering the commissioned ME4/2s, has started. The ME1/1 chambers in the YE+1 nose have meanwhile been re-installed after a total refit in SX5 of the on-chamber and on-disk electronics and power-supply systems. Commissioning will start as soon as cabling and cooling connections are finished. In parallel the maintenance and re-commissioning of CSCs and RPCs in other layers of the positive endcap is ongoing with frequent changes of disk configuration allowing access to different layers.

Other ongoing work is largely independent of the configuration. For instance, on both “blockhouse” platforms the vacuum group is working. At –z, a new gas injection station is under construction allowing faster venting and pumping of the beam pipe in the future, while at +z the corresponding system and also the vacuum pumps are undergoing major refurbishment. The C6F14 cooling plants in UXC are currently being insulated in preparation for Tracker operation at –20 ºC.

Areas of concern

Throughout the entire reporting period intensive work continued inside the vacuum tank at both ends to improve the sealing of the Tracker bulkhead and patch panel regions, deliver dry gas and monitor the humidity and temperature of the environment. A first test in November showed that the cooling plant can deliver cold C6F14, and that the bulkhead environment dew point is well below the lowest likely operating temperature. However, after reverting to room temperature, water was found to be leaking in many places from the insulated pipe-bundles between the cooling distribution plant and the Tracker. Inspection revealed damage in many places to the aluminium vapour barrier of the adhesive foam tape wrapping the insulated pipe bundles. This allows cavern air with a dew point of several tens of degrees to reach the –20 ºC fluorocarbon lines. Ice formation is inevitable and cumulative, also leading to melt-water leaking out of the bundles when cooling is stopped. The potential damage from ice (particularly) and also from melt-water is not tolerable. In reaction to this a very careful inspection and repair campaign has to be planned for January and February before the master-cooling test with the exposed vac tank.

On 15 November, one of two galvanic feed-throughs for heating cables on the outside of ES (–z) was found severely damaged by overheating or arcing. A Technical Incident Panel was held on 25 November and quickly arrived at preliminary conclusions. It rapidly became clear that a repair in situ would be neither feasible nor a sufficient action. Investigations in preparation for the TIP revealed that filter capacitors built into these feed-throughs are not rated for the voltages expected to be used as soon as the ES runs at colder temperatures; moreover the connector pins are not gold-plated. Despite several complicating factors, the most probable explanation of the observed damage is a recent application of higher voltage than previously applied. A clear recommendation was to remove both ES– and ES+ for examination and, at a minimum, repair and exchange of the affected feed-throughs. Space has been prepared in haste in the radiation workshop of SX5. ES+ was removed from the EE on 29 November and on 4 December it was transported to the surface. A first inspection showed no sign of damage on the inside of the feed-through and nominal heating wire resistivity. More detailed tests and checks are ongoing but currently it appears that exchanging the feed-through will solve the problem. The removal to the surface of ES– is foreseen for Monday 9 December.

The construction contractor for the new 43.4 mm (i/d) vacuum chamber (beampipe) for the Phase1 Pixel Tracker has continued to experience problems in completing the final weld, probably related to the new challenge of creating a Beryllium chamber with conical end-pieces. Another attempt in late November produced a mechanically good weld which was unfortunately porous. It was agreed to make one last attempt (with the existing –z conical section and its now very short cylindrical interface) at cutting out the weld and welding in an insert. Meanwhile the manufacturer also agreed to start precautionary pre-machining of a new end-cone with a long cylindrical interface. Fortunately, the latest weld seems mechanically good and has passed initial leak testing. We eagerly await final results. There is still two months’ contingency between the currently estimated “ready for installation” date and the planned installation start.

Planning for the LS1 endgame

The top-priority objective of LS1 is preparing the Tracker to operate at –20 ºC. To address the concerns about vapour barrier damage (above), the configuration of the detector with vac tank exposed for the “master test” of the cooling will be prolonged longer than originally foreseen. This configuration must now be used in parallel for work on the YB+1 and YB+2 and on the positive face of YB0, along with preparatory work for ME1/1 and RE4 (re-)installation at the –z end.

Completing the re-installation and commissioning of ME1/1 at both ends and the construction of the second YE4 (–z), complete with pushback mechanism, are also mandatory LS1 tasks before closure of the detector can start. After the “master cold-test” is over and a feasible Tracker operating temperature determined, the mirror image of today’s logistic configuration has then to be established, but kept for as short a time as reasonably possible before beginning the sequence of closing the detector. This begins with the installation of the beam pipe in early June 2014 and concludes with the full-field magnet test in November 2014. This magnet test will be the ultimate check on the success of sealing the Tracker for colder operation. The handling of this endgame will be difficult and might require some compromises in what can be safely achieved. Technical Coordination plans another LS1 workshop at the end of February to review the status and work out the strategy and schedule options in detail. The very constructive daily interaction with the sub-system field coordinators (and with Run Coordination as commissioning activity builds up) will continue to be vital to translate this into a continually optimised and time-efficient day-to-day planning.

Phase 2 upgrade studies

Priority must be given to successful completion of the LS1 project and the oversight of the remaining approved Phase 1 detector projects (including EYETS 2016–’17 and LS2 planning). However, Technical Coordination intends to pursue, during 2014, a limited set of objectives in support of the work to develop a Technical Proposal for Phase 2. These will focus on the continuing development of a schedule for LS3 not exceeding 30 months, the identification of tasks (mostly infrastructure modifications and end-of-lifetime replacements in preparation for HL-LHC) which can be carried out pre-emptively in LS2, integration and tooling studies for the revision of the YE1 nose (EE/HE/ME1) and (in close consultation with the BRIL project) radiation simulations of particle fluences, beam induced backgrounds, dose rates and activation levels. These simulations are essential to convergence upon an effective Phase 2 design. Associated with these technical objectives, there are serious challenges ahead in developing a viable resource model for common activities over the next decade.

by Austin Ball, Wolfram Zeuner